OpenStudy (vera_ewing):

The equation of a circle is (x + 7)2 + (y + 2)2 = 49. Determine the length of the diameter.

OpenStudy (whpalmer4):

formula for a circle with radius \(r\) and center \((h,k)\) is \[(x-h)^2+(y-k)^2 = r^2\]formula for a diameter \(d\) given a radius \(r\) is \[d = 2r\]

OpenStudy (vera_ewing):

i got 49 is that right? @whpalmer4

OpenStudy (vera_ewing):

@whpalmer4 the choices are 7 9 49 14

OpenStudy (whpalmer4):

Let's take baby steps. Using my formula, what do you find for the center of the circle?

OpenStudy (vera_ewing):

what would i plug into the formula?

OpenStudy (whpalmer4):

\[(x-h)^2 + (y-k)^2 = r^2\]\[(x+7)^2+(y+2)^2 = 49\] what are the values of \(h,k,r\) that make those two equations equal?

OpenStudy (vera_ewing):

7, 2, and 49

OpenStudy (whpalmer4):

I'm sorry, that's 3 strikes...try again?

OpenStudy (vera_ewing):

??? im so confused omg

OpenStudy (whpalmer4):

Hint: do you think \((x-7) = (x+7)\)?

OpenStudy (vera_ewing):

no

OpenStudy (whpalmer4):

Your answer suggests otherwise!

OpenStudy (vera_ewing):

-7, -2, and 49?

OpenStudy (whpalmer4):

now you have 2 of the 3 correct :-) does 49*49 = 49?

OpenStudy (vera_ewing):

2408 is what 49*49=

OpenStudy (vera_ewing):

-7, -2, and 2408

OpenStudy (whpalmer4):

no, 49*49 = 2401 , and that's not 49. for the two equations to be equal, \(r^2 = 49\), right?

OpenStudy (vera_ewing):

yeah

OpenStudy (whpalmer4):

okay, so if \[r^2=49\]what does \(r=\)

OpenStudy (vera_ewing):

r=7

OpenStudy (whpalmer4):

good. if \(r=7\), what does \(d=\)

OpenStudy (vera_ewing):

d=14?

OpenStudy (whpalmer4):

as an old math teacher of mine used to say when someone answered with uncertainty in their voice, "is that an answer, or a prayer?" :-)

OpenStudy (vera_ewing):

d=14!

OpenStudy (whpalmer4):

there shouldn't be any need for uncertainty as to the value of \(d = 2r\) if you know that \(r = 7\) that's better!

OpenStudy (vera_ewing):

thank you so much! can you help with one more? I'll ask a new question :)

OpenStudy (whpalmer4):

so, this sort of pattern-matching is a very useful and important thing to grasp. let's do another one as practice: \[(x+1)^2 + (y-3)^2 = 25\] what is the center of that circle, and what is the radius? again, the formula for the circle with center at \((h,k)\) and radius \(r\) is: \[(x-h)^2 + (y-k)^2 = r^2\]

OpenStudy (vera_ewing):

ok so the center is -1, the radius is 3

OpenStudy (whpalmer4):

:-( isn't the center a point with 2 coordinates? and does 3*3 = 25?

OpenStudy (vera_ewing):

ugh im so confused again! im supposed to multiply?

OpenStudy (whpalmer4):

let's go through it again: \[(x+1) = (x-h)\]\[x+1 = x-h\]subtract \(x\) from both sides \[+1 = -h\]\[h = -1\] \[(y-3) = (y-k)\]\[y-3=y-k\]\[-3=-k\]\[k=3\] so the center is at \((h,k) = (-1,3)\)

OpenStudy (whpalmer4):

looking at the right side of the equation, \[r^2 = 25\]\[r =\]

OpenStudy (vera_ewing):

yeah ok i get that. i don't know why you asked if 3*3 is 25 though??

OpenStudy (vera_ewing):

r=5

OpenStudy (whpalmer4):

you said the radius was 3! I'm saying you couldn't possibly have checked your answer for reasonableness, unless you think that 3*3 = 25, because the value of the radius ^2 has to equal 25

OpenStudy (vera_ewing):

oh ok

OpenStudy (whpalmer4):

you had another problem?

OpenStudy (vera_ewing):

yeah i"ll ask it right now. it's a question from a practice test.